Old MacDonald wants to build a rectangular pen for his animals. He has 40 m of fencing and he plans to use the barn as one side for his pen. What dimensions will give him a maximum area?
Answer by checkley71(8403)

I APOLOGIZE FOR THE INCORRECT ANSWER.
IT SEEMS THAT A 10 BY 20 BY 10 RECTANGULAR PEN GIVES AN AREA OF 10*20=200 M^2 WHICH IS THE MAX AREA. I OVERLOOKED THE WORD RECTANGULAR BECAUSE A SQUARE IS A SPECIAL RECTANGLE.
-----------------------------------------------------------------------------
A SQUARE IS THE MAXIMUM AREA WITH A MINIMUM PERIMETER.
HOWEVER YOU ONLY NEED 3 SIDES OF FENCE.
THUS YOU NEED TO DIVIDE YOUR FENCING BY 3 TO GET THE SQUARE.
40/3=13.333 M IS THE LENGTH OF THE EACH SIDE.
THIS GIVES AN AREA OF:
13.333*13.333=177.769 M^2.